$11,000 is invested for 9 years with an APR of 4% and quarterly compounding, what is the balance after 9 years?
Answer by rfer(16322)

FV=11000(1+0.04/4)^4*9
FV=11000(1.01)^36
FV=11000*1.43
FV=$15,730